The market demand and supply equations for theme park in a city are given by P = 30 – 0.005QD and P = 10 + 0.005QS, where P is the price in dollars and QD is the quantity of theme-park tickets demanded and QS is the quantity of theme park ticket supplied.


Explain the implications of the welfare of consumers, producers and the society when the price of theme park ticket is fixed at $15. Support your answers with a graph of the theme park tickets market.


1
Expert's answer
2022-02-07T08:33:27-0500

In equilibrium Pd = Ps, so:

"30 \u2013 0.005Q = 10 + 0.005Q,"

0.01Q = 20,

Q = 2,000 units,

"P = 10 + 0.005*2,000 = 20."

If the price of theme park ticket is fixed at $15, which is below equilibrium, then the deficit of theme-park tickets and the deadweight-loss will occur.